OpenStudy (anonymous):

simplify the square root of 8 + the square root of 50

OpenStudy (anonymous):

\[2\sqrt{2}+5\sqrt{2}=5\sqrt{2}\]

OpenStudy (dumbcow):

sqrt(8) = sqrt(4*2) = sqrt(4)*sqrt(2) = 2sqrt(2) sqrt(50) = sqrt(25*2) = sqrt(25)*sqrt(2) = 5sqrt(2) add like terms 2+5 = 7 = 7sqrt(2)

OpenStudy (anonymous):

should be \[7\sqrt{2}\] sorry

OpenStudy (anonymous):

\[\sqrt{8}+\sqrt{50}=\sqrt{(2)(4)}+\sqrt{(2)(25)} =\sqrt{2}\sqrt{4}+\sqrt{2}\sqrt{25} =2\sqrt{2}+5\sqrt{2}=\sqrt{2}(2+5)=7\sqrt{2}\]